OpenStudy (anonymous):

The sum of the two digits of a number is 9. If the tens digit is one-half the units digit, what is the number? Let t = the tens digit, u = the units digit, and t + u = 9. Which of the following equations would complete the system?

OpenStudy (anonymous):

answers t=1/2u u=1/2t t-u=1/2

OpenStudy (jdoe0001):

\(\bf \text{If the }\color{blue}{\text{tens digit is one-half the units digit,}}\\ \text{ what is the number?} \)

OpenStudy (jdoe0001):

the number is "tu" where "t" is half of "u"

OpenStudy (anonymous):

the no. is 36